# Privacy-Preserving Federated Learning Scheme with Mitigating Model Poisoning Attacks: Vulnerabilities and Countermeasures

### v1

@@ 4.4 Immediate Fix and Its Ineffectiveness
The main reason for the privacy breaches in ShieldFL is that the local gradients $g_{i}^{(t)}=\{x_{i1},\cdots,x_{il}\}$ and the last-iteration aggregated gradient $g_{\Sigma}^{(t-1)}=\{y_{1},\cdots,y_{l}\}$ (or the poisonous gradient baseline $g_{*}^{(t)}=\{z_{1},\cdots,z_{l}\}$) are masked with the same random noise vector, thereby exposing the correlations between these gradients and resulting in the severe consequence that once one of these gradients is leaked, all other gradients can be directly revealed in clear.
Although different random noise vectors can be used on the local gradients, the last-iteration aggregated gradient, and the poisonous gradient baseline to prevent the aforementioned privacy violations, using the same noise vector is inevitable to maintain the correctness of the secure cosine similarity measurement $\mathbf{SecCos}$. In detail, the secure cosine similarity measurement between two gradient ciphertexts $[\![g_{a}]\!]$ and $[\![g_{b}]\!]$ is calculated by first masking the two gradients with random noise vectors $r$ and $\gamma$ to output $[\![\bar{g}_{a}]\!]=[\![g_{a}+r]\!]$ and $[\![\bar{g}_{b}]\!]=[\![g_{b}+\gamma]\!]$, respectively, then decrypting the masked gradients and calculating the cosine similarity $\overline{\cos}_{ab}=\frac{\bar{g}_{a}\odot\bar{g}_{b}}{|\!|\bar{g}_{a}|\!|\cdot|\!|\bar{g}_{b}|\!|}=\frac{(g_{a}+r)\odot(g_{b}+\gamma)}{|\!|g_{a}+r|\!|\cdot|\!|g_{b}+\gamma|\!|}$. However, $\frac{\bar{g}_{a}\odot\bar{g}_{b}}{|\!|\bar{g}_{a}|\!|\cdot|\!|\bar{g}_{b}|\!|}=\frac{g_{a}\odot g_{b}}{|\!|g_{a}|\!|\cdot|\!|g_{b}|\!|}$ holds only when $r=\gamma$, where $\frac{g_{a}\odot g_{b}}{|\!|g_{a}|\!|\cdot|\!|g_{b}|\!|}$ is the correct computation of the cosine similarity between $g_{a}$ and $g_{b}$; Otherwise, if $r\neq\gamma$, the correctness of the calculation will be breached.
@@ 5.1 Technical Intuition
We empower our enhanced model to resist privacy attacks while ensuring correctness. In ShieldFL’s secure computation method, the server $S_{1}$ employs masking techniques to process the user gradient ciphertexts it receives, then $S_{1}$ and $S_{2}$ interact to compute on the masked ciphertexts, and finally $S_{2}$ partially decrypts the computation results of the masked ciphertexts and sends them back to $S_{1}$. $S_{1}$ obtains the masked computation results and demasks to acquire the final computation plaintext output. While the masking technique is highly efficient, ensuring the correctness of ciphertext computations requires that different ciphertexts be masked with the same random noise, which leads to significant privacy leakage issues.
Our enhanced scheme retains the efficient masking techniques to safeguard the security of interactions between the two servers. However, unlike ShieldFL, each ciphertext in our scheme is masked with randomly selected noise. To ensure correctness of computations on the masked ciphertexts, we categorize operations on ciphertexts into linear (such as addition, subtraction, and scalar multiplication) and non-linear operations (such as multiplication and exponentiation) from the perspective of masking. For linear operations, the computation results of the masked ciphertexts can be correctly demasked; however, non-linear operations introduce errors that are challenging to correct. To address this issue without adding computational burdens, we perform non-linear operations on the original ciphertexts locally at $S_{1}$. After conducting these operations, $S_{1}$ masks the results and sends them to $S_{2}$, which then performs linear operations on the masked ciphertexts and decrypts the results. Consequently, $S_{1}$ can successfully demask $S_{2}$’s decrypted results to obtain the final correct computation output. In our scheme, $S_{1}$ only accesses the final computation results, and $S_{2}$ retains the masked intermediate computation results, thus resisting our privacy attacks.
@@
(1) Ciphertext computation simplicity: In our enhanced scheme, the secure computation process of non-linear operation on ciphertexts–crucial for operations like computing ${\cos}_{ab}={g}_{a}\odot{g}_{b}$ during $\mathbf{SecCos}([\![g_{a}]\!],[\![g_{b}]\!])$–is streamlined. Unlike previous methods that relied on Paillier’s addition homomorphism and complex masking and unmasking techniques, requiring multiple interactions between the two servers, we construct a two-trapdoor FHE that allows for direct non-linear operations on ciphertexts. This direct approach can significantly reduces both the number of computation steps and communication rounds, making our secure computation methods more efficient than those in ShieldFL. For example, our secure computation methods (outlined in Figs. 3 and 4) only require two rounds of interaction between $S_{1}$ and $S_{2}$, whereas the original scheme requires four rounds, reducing the communication burden by (at least) half.
(2) Cryptographic primitive efficiency: We employ an RLWE-based FHE, specifically the CKKS cryptosystem, to construct our two-trapdoor FHE scheme. This choice is more efficient than the Paillier cryptosystem-based two-trapdoor PHE. The RLWE-based system, with its smaller ciphertext modulus requirements, offers lower computational overhead and communication traffic per step. For instance, at a $128$-bit security level, the two-trapdoor PHE demands a ciphertext modulus $N^{2}$ of at least $3072\times 2$ bits, whereas the RLWE-based FHE can operate with a much smaller $q$ of $56$ bits (with $n_{f}=2048$), significantly reducing the computational cost. We provide a detailed performance comparison in Section 6.2.
In summary, our enhanced scheme introduces significant improvements in both efficiency and security in the realm of secure computation. By employing randomly selected noise for masking and categorizing operations into linear and non-linear types, we effectively address the privacy concerns associated with previous methods. The streamlined process for non-linear operations, facilitated by a two-trapdoor FHE based on the RLWE homomorphic encryption scheme, minimizes the complexity and communication overhead typically required. This innovative approach not only ensures the correctness of the computations but also bolsters resistance to privacy attacks, setting a new benchmark for secure computation methods. Ultimately, our advancements position our scheme as a more efficient and secure alternative to existing solutions like ShieldFL.
@@ 5.2 New Secure Computation Methods
  • $\mathbf{FHE.Enc}_{pk}(m)\rightarrow[\![m]\!]$: This algorithm first encodes a float vector $m\in\mathbb{R}^{n_{f}}$ with length $n_{f}$ into an integral coefficient polynomials $m(X)\in\mathcal{R}_{p}$ and then it encrypts $m(X)$ and outputs the ciphertext $[\![m]\!]$. Specifically, the input vector $m$ is first converted to a complex vector $z=(z_{1},\cdots,z_{n_{f}/2})\in\mathbb{C}^{n_{f}/2}$, in which every two elements in $m$ form the real and imaginary parts of a complex number. Then, the complex vector $z$ is encoded as $m(X)=\lfloor\Delta\cdot\phi^{-1}(z)\rceil$, where $\Delta>1$ is a scaling factor and $\phi:R[X]/(X^{n_{f}}+1)\rightarrow\mathbb{C}^{n_{f}/2}$ is a ring isomorphism. Finally, the ciphertext $[\![m]\!]=(m+u\cdot pk_{0}+e_{0}\bmod q,u\cdot pk_{1}+e_{1}\bmod q)\triangleq(c_{0},c_{1})\in\mathcal{R}_{q}^{2}$ is output, where $u\leftarrow\chi_{s}$ and $e_{0},e_{1}\leftarrow\chi_{e}$.
@@
  • $\mathbf{FHE.FullDec}([\![m]\!],[m]_{1},[m]_{2})\rightarrow m$: Given a ciphertext $[\![m]\!]=(c_{0},c_{1})$ and its partial decryption result pair $([m]_{1},[m]_{2})$, the algorithm first fully decrypts the ciphertext $[\![m]\!]$ as a plaintext polynomial $m(X)=c_{0}+[m]_{1}+[m]_{2}\bmod p\in\mathcal{R}_{p}$, then it decodes $m(X)$ by first computing the complex vector $z=\Delta^{-1}\cdot\phi(m(X))\in\mathbb{C}^{n_{f}/2}$ and then converting $z$ to a float vector $m\in\mathbb{R}^{n_{f}}$. Finally, the algorithm outputs $m$.
@@ 5.3 The Enhanced Scheme
  1. 1.
    Setup: In this process, $KC$ generates public parameters/keys for the scheme and $S_{1}$ generates an initial model for the users. Specifically, given a security parameter $\varepsilon^{\prime}$, $KC$ calls $\mathbf{FHE.Setup}$ and $\mathbf{FHE.KeyGen}$ to generate the public parameter $PP$ and a tuple of key $(sk,pk,evk)$. Then, $KC$ calls $\mathbf{FHE.KeySplit}(sk)$ with $n+1$ times to generate $n+1$ pairs of secret key shares $(sk_{1},sk_{2})$ and $\{(sk_{s_{i}},sk_{u_{i}})\}_{i\in[1,n]}$. Finally, $KC$ broadcasts $\{PP,pk,evk\}$, distributes $sk_{1},\{sk_{s_{i}}\}_{i\in[1,n]}$ to $S_{1}$, distributes $sk_{2}$ to $S_{2}$, and distributes $sk_{u_{i}}$ to $U_{i}$. $S_{1}$ selects a random initial model $W^{(0)}$ and sends $W^{(0)}$ to $U_{i}$.
  2. 2.
    Local Training: In each training iteration, each user trains its local model to obtain a local gradient, encrypts the local gradient, and sends the encrypted gradient to $S_{1}$. Specifically, in the $t$th training iteration, each user $U_{i}$ trains its local model $W^{(t)}$ and flats its local gradient as a gradient vector $g_{i}^{(t)}\in\mathbb{R}^{l}$ whose length is $l$. The gradient vector is then normalized and uniformly divided into $\tau=\lceil l/n_{f}\rceil$ vectors $\{g_{i,j}^{(t)}\in\mathbb{R}^{n_{f}}\}_{j\in[1,\tau]}$ with length $n_{f}$. The elements 0s are padded at the end of $g_{i}^{(t)}$ when less than $n_{f}$ elements are left. Then, $U_{i}$ encrypts each vector $g_{i,j}^{(t)}$ as $[\![g_{i,j}^{(t)}]\!]$ by calling $\mathbf{FHE.Enc}$. Finally, the beginning user $U_{i}$ sends its encrypted gradient $[\![g_{i}^{(t)}]\!]=\{[\![g_{i,1}^{(t)}]\!],\cdots,[\![g_{i,\tau}^{(t)}]\!]\}$ to $S_{1}$ and the malicious user $U_{i}\in U^{*}$ may launch poisoning attacks and then send its encrypted poisonous gradients $[\![g_{i*}^{(t)}]\!]=\{[\![g_{i*,1}^{(t)}]\!],\cdots,[\![g_{i*,\tau}^{(t)}]\!]\}$ to $S_{1}$.
  3. 3.
    Privacy-Preserving Defense: This process mainly contains the following three implementations.
    • Normalization judgment. $S_{1}$ determines whether its received local gradient ciphertexts $[\![g_{i}^{(t)}]\!]=$$\{[\![g_{i,1}^{(t)}]\!],\cdots,[\![g_{i,\tau}^{(t)}]\!]\}$ are normalized by calling $\mathbf{ESecJudge}([\![g_{i}^{(t)}]\!])$.
    • Poisonous gradient baseline finding. First, $S_{1}$ calculates the cosine similarity between the last-iteration aggregated gradient ciphertext $[\![g_{\Sigma}^{(t-1)}]\!]$ and each participant’s local gradient ciphertext $[\![g_{i}^{(t)}]\!]$ as
      $\cos_{i\Sigma}\leftarrow\mathbf{ESecCos}([\![g_{i}^{(t)}]\!],[\![g_{\Sigma}^{(t-1)}]\!]),\textrm{for\ }i\in[1,n].$
      Then, $S_{1}$ chooses the poisonous baseline $[\![g_{*}^{(t)}]\!]$ from the local gradient ciphertexts $\{[\![g_{i}^{(t)}]\!]\}_{i\in[1,n]}$, whose cosine similarity $\cos_{*\Sigma}$ is the lowest one among $\{\cos_{i\Sigma}\}_{i\in[1,n]}$.
    • Byzantine-tolerance aggregation. To motivate users to engage in benign training and accurately identify poisonous gradients, we construct the Byzantine-tolerance aggregation method as demonstrated below.
      • $S_{1}$ initiates the credit score $cs_{i}=100\%$ for the user $i(i\in[1,N])$.
      • $S_{1}$ computes the cosine similarity between the poisonous baseline $[\![g_{*}^{(t)}]\!]$ and the local gradient ciphertext $[\![g_{i}^{(t)}]\!]$ as $\cos_{i*}\leftarrow\mathbf{ESecCos}([\![g_{i}^{(t)}]\!],[\![g_{*}^{(t)}]\!]),\textrm{for\ }i\in[1,n].$ The smaller $\cos_{i*}$ is, the more likely that $[\![g_{i}^{(t)}]\!]$ is a poisonous gradient. Note that $\cos_{i*}\in\mathbb{Z}_{p}$.
      • $S_{1}$ sets the confidence $co_{i}\leftarrow(p-\cos_{i*})\times cs_{i}$ for the gradient $g_{i}^{(t)},i\in(1,N)$ instead of eliminating outliers, to adapt to non-IID datasets among users.
      • $S_{1}$ recomputes all users’ credit scores $cs_{i}=cs_{i}\times\lfloor\frac{co_{i}}{co}\rceil,i\in(1,N)$ for the next iteration, where $co\leftarrow\sum_{j=1}^{n}co_{j}$. The lower the confidence, the more likely the user is malicious, and thus the user’s credit score needs to be reduced accordingly, and on the contrary, it needs to be increased. The credit score contains all history confidences of the user and thus incentives the user to engage in benign training in the whole training process.
      • $S_{1}$ calculates $co_{i}\leftarrow\lfloor\dfrac{co_{i}}{co}\cdot p\rceil$ and weights the local gradients with their respective confidences and then aggregates them to generate a global gradient $[\![g_{\Sigma}^{(t)}]\!]\leftarrow\sum_{i=1}^{n}[\![co_{i}]\!][\![g_{i}^{(t)}]\!]$, where $g_{\Sigma}^{(t)}=\sum_{i=1}^{n}co_{i}g_{i}^{(t)}$.
  4. 4.
    Model Update. $S_{1}$ sends $co,[\![g_{\Sigma}^{(t)}]\!]$ and $[g_{\Sigma}^{(t)}]_{s_{i}}\leftarrow\mathbf{FHE.PartDec}_{sk_{s_{i}}}([\![g_{\Sigma}^{(t)}]\!])$ to $U_{i}$. Then $U_{i}$ calculates the plaintext aggregated gradient $g_{\Sigma}^{(t)}$ by calling $\mathbf{FHE.PartDec}$ and $\mathbf{FHE.FullDec}$ with $sk_{u_{i}}$. Finally, $U_{i}$ updates its local model as $W^{(t+1)}\leftarrow W^{(t)}-co^{-1}\eta^{(t)}g_{\Sigma}^{(t)}$, where $\eta^{(t)}$ is the learning rate.
TABLE III: Computational costs.
Schemes Setup Local Training Privacy-Preserving Defense Model Update Total
$O(\varepsilon+(n+1)\cdot ks$ $O(nlT\cdot train$ $O(4nlT(N\cdot rnd+penc$ $O(3nlT\cdot pdec$ $O(\varepsilon+(n+1)\cdot ks$
ShieldFL $+l\cdot penc$ $+nlT\cdot penc))$ $+9\cdot pdec+padd$ $+nlT\cdot update)$ $+nlT(6N^{4}\cdot mul+40N^{3}\cdot add$
$+(2n+1)l\cdot pdec)$ $+N^{2}\cdot pmul))$ $+8N^{3}\cdot rnd+train+update)$
$O(\varepsilon^{\prime}+(n+1)n_{f}\cdot ks$ $O(nlT\cdot train$ $O(n\tau T(3q\cdot rnd+3fenc$ $O(3n\tau T\cdot fdec$ $O(\varepsilon^{\prime}+(n+1)n_{f}\cdot ks$
Our scheme $+\tau\cdot fenc$ $+n\tau T\cdot fenc))$ $+9fdec+5fadd$ $+nlT\cdot update)$ $+n\tau T(35n_{f}q\cdot mul+59n_{f}q\cdot add$
@@
TABLE IV: Communication costs.
Schemes Setup Local Training Privacy-Preserving Defense Model Update Total
ShieldFL $O((2n+2)\varepsilon+2nlC_{p})$ $O(nlTC_{p})$ $O(nT(10l+9)C_{p})$ $O(2nlTC_{p})$ $O(2n\varepsilon+13nlTN^{2})$
Our scheme $O((2n+2)n_{f}\varepsilon^{\prime}+2n\tau C_{f})$ $O(n\tau TC_{f})$ $O(nT(2\tau+4)C_{f})$ $O(2n\tau TC_{f})$ $O(2nn_{f}\varepsilon^{\prime}+5n\tau Tq)$

### v2

@@ 4.2 Immediate Fix and Its Ineffectiveness
The main reason for the privacy breaches in ShieldFL is that the local gradients $g_{i}^{(t)}$ and the last-iteration aggregated gradient $g_{\Sigma}^{(t-1)}$ (or the poisonous gradient baseline $g_{*}^{(t)}$) are masked with the same random noise vector, thereby exposing the correlations between these gradients and resulting in the severe consequence that once one of these gradients is leaked, all other gradients can be directly revealed in clear.
Although different random noise vectors can be used to prevent the aforementioned privacy violations, using the same noise vector is inevitable to maintain the correctness of $\mathbf{SecCos}$. In detail, the secure cosine similarity measurement between two gradient ciphertexts $[\![g_{a}]\!]$ and $[\![g_{b}]\!]$ is calculated by first masking the two gradients with random noise vectors $r$ and $\gamma$ to output $[\![\bar{g}_{a}]\!]=[\![g_{a}+r]\!]$ and $[\![\bar{g}_{b}]\!]=[\![g_{b}+\gamma]\!]$, respectively, then decrypting the masked gradients and calculating the cosine similarity $\overline{\cos}_{ab}=\frac{\bar{g}_{a}\odot\bar{g}_{b}}{|\!|\bar{g}_{a}|\!|\cdot|\!|\bar{g}_{b}|\!|}=\frac{(g_{a}+r)\odot(g_{b}+\gamma)}{|\!|g_{a}+r|\!|\cdot|\!|g_{b}+\gamma|\!|}$. However, $\frac{\bar{g}_{a}\odot\bar{g}_{b}}{|\!|\bar{g}_{a}|\!|\cdot|\!|\bar{g}_{b}|\!|}=\frac{g_{a}\odot g_{b}}{|\!|g_{a}|\!|\cdot|\!|g_{b}|\!|}$ holds only when $r=\gamma$, where $\frac{g_{a}\odot g_{b}}{|\!|g_{a}|\!|\cdot|\!|g_{b}|\!|}$ is the correct computation of the cosine similarity between $g_{a}$ and $g_{b}$; Otherwise, if $r\neq\gamma$, the correctness will be breached.
@@ 5.1 Technical Intuition
We empower our enhanced scheme to resist privacy attacks while ensuring computation correctness. In ShieldFL’s secure computation methods, the server $S_{1}$ applies a masking technique to the received user gradient ciphertexts. Then $S_{1}$ and $S_{2}$ interact to compute on the masked ciphertexts, and finally $S_{2}$ partially decrypts the computation results of the masked ciphertexts and sends them back to $S_{1}$. $S_{1}$ obtains the masked computation results and demasks to acquire the final computation plaintext output. Although the masking technique is highly efficient, ensuring the correctness of computation requires that all ciphertexts be masked using the same random noise. This introduces strong correlations among masked values: if any plaintext is leaked, all correlated gradients can be inferred, thereby violating privacy. A straightforward solution, using independent noise for each ciphertext, would eliminate these correlations but compromise computation correctness.
To simultaneously ensure both security and correctness, we fundamentally redesign the secure computation paradigm in our enhanced scheme. Specifically, we introduce a novel masking protocol build upon a newly constructed two-trapdoor FHE scheme. Instead of masking the inputs, we mask the computation results using different random vectors. This eliminates the need for shared noise and preserves correctness. To handle ciphertext operations under masking, we classify them into linear (e.g., addition, subtraction, scalar multiplication) and non-linear (e.g., multiplication, exponentiation) categories. Linear operations on masked ciphertexts yield correctly demaskable results. In contrast, non-linear operations introduce complex noise that is difficult to eliminate. To mitigate this without adding computational overhead, we perform non-linear operations directly on the original ciphertexts at $S_{1}$. The results are then masked and sent to $S_{2}$, which performs the remaining linear operations and partial decryption. Finally, $S_{1}$ removes the mask to obtain the correct plaintext output. In this design, $S_{1}$ only observes the final computation results, while $S_{2}$ only accesses masked intermediate values. This separation effectively defends against the proposed privacy attacks.
@@
(1) Ciphertext computation simplicity: In our enhanced scheme, the secure computation process of non-linear operations on ciphertexts – crucial for operations like computing ${\cos}_{ab}=g_{a}\odot g_{b}$ during $\mathbf{SecCos}([\![g_{a}]\!],[\![g_{b}]\!])$ – is streamlined. Unlike previous methods that relied on Paillier’s addition homomorphism and complex masking/unmasking techniques, requiring multiple interactions between the two servers, we construct a two-trapdoor FHE that allows for direct non-linear operations on ciphertexts. This direct approach significantly reduces both the number of computation steps and the communication rounds, making our secure computation methods more efficient than those in ShieldFL. For example, our secure computation methods (outlined in Figs. 3 and 4) only require two rounds of interaction between $S_{1}$ and $S_{2}$, whereas the original scheme requires four rounds, reducing the communication burden by (at least) half.
(2) Cryptographic primitive efficiency: We employ an RLWE-based FHE, specifically the CKKS cryptosystem, to construct our two-trapdoor FHE scheme. This choice is more efficient than the Paillier cryptosystem-based two-trapdoor PHE. The RLWE-based system, with its smaller ciphertext modulus requirements, offers lower computational overhead and communication traffic per step. For instance, at a 128128128-bit security level, the two-trapdoor PHE demands a ciphertext modulus $N^{2}$ of at least 3072×23072$\times$ 23072 × 2 bits, whereas the RLWE-based FHE can operate with a much smaller $q$ of 565656 bits (with $n_{f}=2048$), significantly reducing the computational cost. We provide a detailed performance comparison in Section 6.3.
(3) No additional trade-offs: Importantly, these improvements do not introduce extra limitations. Our enhanced scheme retains the same flexibility and scalability as ShieldFL because both adopt a two-trapdoor architecture and the same key distribution mechanism. Although CKKS introduces negligible approximation error due to its encoding and encryption, such error is naturally tolerated by stochastic machine learning models and does not impact convergence or accuracy, as verified by our experimental results in Section LABEL:sec:experiment.
In summary, our enhanced scheme introduces significant improvements in both efficiency and security without sacrificing flexibility, accuracy, or scalability. By employing randomly selected noise for masking and categorizing operations into linear and non-linear types, we effectively address the privacy concerns associated with previous methods. The streamlined process for non-linear operations, facilitated by a two-trapdoor FHE based on the RLWE homomorphic encryption scheme, minimizes the complexity and communication overhead typically required. This innovative approach not only ensures the correctness of the computations but also bolsters resistance to privacy attacks, setting a new benchmark for secure computation methods. Ultimately, our advancements position our scheme as a more efficient and secure alternative to existing solutions like ShieldFL.
@@ 5.2 New Secure Computation Methods
[…]
@@
[…]
@@ 5.3 Construction of Our PBFL Scheme
2) Local Training: In each training iteration, each user trains its local model to obtain a local gradient, encrypts the local gradient, and sends the encrypted gradient to $S_{1}$. Specifically, in the $t$th training iteration, each user $U_{i}$ trains its local model $W^{(t)}$ and flats its local gradient as a gradient vector $g_{i}^{(t)}\in\mathbb{R}^{l}$ whose length is $l$. The gradient vector is then normalized and uniformly divided into $\tau=\lceil l/n_{f}\rceil$ vectors $\{g_{i,j}^{(t)}\in\mathbb{R}^{n_{f}}\}_{j\in[1,\tau]}$ with length $n_{f}$. The elements 0s are padded at the end of $g_{i}^{(t)}$ when less than $n_{f}$ elements are left. Then, $U_{i}$ encrypts each vector $g_{i,j}^{(t)}$ as $[\![g_{i,j}^{(t)}]\!]$ by calling $\mathbf{FHE.Enc}$. Finally, the beginning user $U_{i}$ sends its encrypted gradient $[\![g_{i}^{(t)}]\!]=\{[\![g_{i,1}^{(t)}]\!],\cdots,[\![g_{i,\tau}^{(t)}]\!]\}$ to $S_{1}$ and the malicious user $U_{i}\in U^{*}$ may launch poisoning attacks and then send its encrypted poisonous gradients $[\![g_{i*}^{(t)}]\!]=\{[\![g_{i*,1}^{(t)}]\!],\cdots,[\![g_{i*,\tau}^{(t)}]\!]\}$ to $S_{1}$.
3) Privacy-Preserving Defense: This process mainly contains the following three implementations.
  • Normalization Judgment. The server $S_{1}$ determines whether the received encrypted local gradients $[\![g_{i}^{(t)}]\!]=\{[\![g_{i,1}^{(t)}]\!],\cdots,[\![g_{i,\tau}^{(t)}]\!]\}$ are regularized by invoking the verification function $\mathbf{ESecJudge}([\![g_{i}^{(t)}]\!])$. Only those passing the verification are considered trusted and included in the index set $\mathcal{I}_{\text{trusted}}$.
  • Poisonous Gradient Baseline Identification. For all trusted gradients $\{[\![g_{i}^{(t)}]\!]\}_{i\in\mathcal{I}_{\text{trusted}}}$, the server computes the cosine similarity with the previously aggregated global gradient $[\![g_{\Sigma}^{(t-1)}]\!]$ as $\cos_{i\Sigma}\leftarrow\mathbf{ESecCos}([\![g_{i}^{(t)}]\!],[\![g_{\Sigma}^{(t-1)}]\!]),\quad\text{for }i\in\mathcal{I}_{\text{trusted}}.$ The gradient having the lowest similarity score is selected as the poisonous baseline $[\![g_{*}^{(t)}]\!]$.
  • Byzantine-Tolerant Aggregation. To defend against adversarial behaviors while incentivizing honest participants, the server constructs a dual-scoring aggregation strategy using both confidence and credit scores.
    • Confidence Computation. For each user $i\in\mathcal{I}_{\text{trusted}}$, the cosine similarity between their local gradient and the poisonous baseline is computed as
      $\cos_{i*}\leftarrow-\ \mathbf{ESecCos}([\![g_{i}^{(t)}]\!],[\![g_{*}^{(t)}]\!]).$
      These similarity values are then normalized to obtain confidence scores:
      $co_{i}\leftarrow\frac{\exp(\cos_{i*}-\max_{j}\cos_{j*})}{\sum_{j\in\mathcal{I}_{\text{trusted}}}\exp(\cos_{j*}-\max_{k}\cos_{k*})}.$
    • Credit Score Update. Each participant¡¯s credit score $cs_{i}$ is updated with exponential smoothing
      $cs_{i}\leftarrow\alpha\cdot cs_{i}+(1-\alpha)\cdot co_{i},$
      where $\alpha\in[0.7,0.95]$ is a decay coefficient to balance historical consistency and current behavior.
    • Client Aggregation Weights. The clients’ aggregation weights are computed based on the product of credit and confidence scores
      $w_{i}\leftarrow\frac{cs_{i}\cdot co_{i}}{\sum_{j\in\mathcal{I}_{\text{trusted}}}cs_{j}\cdot co_{j}}.$
    • Adaptive Client Filtering. To further enhance robustness against adversarial manipulation and reduce the influence of anomalous or overconfident updates, the server performs an optional three-stage client filtering strategy before final aggregation:
      • (i)
        Adaptive Weight Mixture. The server first interpolates between uniform weighting and the learned robust weights based on the current round index $t$, using a decaying mixture coefficient:
        $\lambda\leftarrow\max\left(0,1-\frac{t-T_{\text{warmup}}}{T_{\text{total}}-T_{\text{warmup}}}\right),$
        $\tilde{w}_{i}\leftarrow\lambda\cdot\frac{1}{n}+(1-\lambda)\cdot w_{i},$
        where $\tilde{w}_{i}$ denotes the adjusted aggregation weight for client $i$, and $n$ is the total number of clients. This allows the server to rely more on uniform weighting in early rounds and gradually transition to confidence-driven weighting as training progresses. Clients with $\tilde{w}_{i}<\theta$ (e.g., $\theta=0.05$) are excluded from the aggregation.
      • (ii)
[…]
      • (iii)
        Credit Penalty for Filtered Clients. To enhance long-term robustness, clients filtered out during either the Normalization Judgment or Outlier Suppression stages are penalized by applying differentiated scaling factors to their credit scores. Specifically, a multiplicative decay $\gamma_{1}\in[0,1]$ is applied to clients excluded by normalization, while an inflation factor $\gamma_{2}>1$ is used for clients removed during outlier suppression. This adjustment modifies each client’s trust trajectory in future rounds accordingly.
    Finally, the weighted aggregation of encrypted gradients is computed as
    $[\![g_{\Sigma}^{(t)}]\!]\leftarrow\sum_{i\in\mathcal{I}_{\text{select}}}\frac{w_{i}}{\sum_{j\in\mathcal{I}_{\text{select}}}w_{j}}\cdot[\![g_{i}^{(t)}]\!].$
4) Model Update: The server $S_{1}$ partially decrypts and sends the share $[g_{\Sigma}^{(t)}]_{s_{i}}\leftarrow\mathbf{FHE.PartDec}_{sk_{s_{i}}}([\![g_{\Sigma}^{(t)}]\!])$ to each participant $U_{i}$. Each user completes decryption using its private key $sk_{u_{i}}$
$g_{\Sigma}^{(t)}\leftarrow\mathbf{FHE.FullDec}(\{[g_{\Sigma}^{(t)}]_{s_{i}}\},sk_{u_{i}}),$
and updates its local model using
$W^{(t+1)}\leftarrow W^{(t)}-\eta^{(t)}\cdot g_{\Sigma}^{(t)},$
where $\eta^{(t)}$ is the learning rate of round $t$.
Our proposed privacy-preserving defense is designed to be robust against both conventional data poisoning attacks (as examined in [26]) and more sophisticated adaptive model poisoning attacks, including the AGR-tailored and AGR-agnostic strategies introduced in [42]. To counter these evolving threats, our defense integrates several key mechanisms that jointly enhance resilience:
  • Dual-Scoring Trust Assessment. A two-tier scoring system is employed, combining per-round cosine-based confidence scores and long-term credit scores. This joint quantification reflects both short-term gradient consistency and historical client reliability, effectively preventing adaptive adversaries from persistently imitating benign behaviors to evade detection.
  • Exponential Credit Score Smoothing. Credit scores are updated via exponential smoothing with decay factor $\alpha\in[0.7,0.95]$, enabling the system to gradually penalize clients with recurrent suspicious updates while remaining tolerant to occasional noise in benign gradients.
  • Soft Confidence Normalization. Instead of applying hard thresholding, we utilize softmax-based normalization of negative cosine similarities to compute smooth confidence scores. This soft treatment maintains sensitivity to fine-grained deviations and provides robustness against gradient-level perturbations crafted by adaptive attackers.
  • Adaptive Weight Mixing and Thresholding. An adaptive mixing scheme transitions from uniform weighting to fully trust-weighted aggregation as training progresses, enhancing early-round stability. Clients with extremely low trust weights (below threshold $\theta$) are excluded from aggregation to reduce stealth attack surfaces.
  • Outlier Suppression via Drop Detection. Our outlier suppression step explicitly targets highly confident yet anomalously extreme updates, which are likely crafted to exploit the statistical patterns of benign gradients for malicious purposes. This mechanism is particularly critical for defending against adaptive adversaries, such as those employing AGR-tailored strategies. In such attacks, malicious gradients are deliberately constructed based on the distributional characteristics of benign updates, typically by adding a scaled deviation vector to their mean. These adversarial updates often evade baseline filters due to their close proximity to the benign gradient space. Furthermore, they may achieve abnormally high confidence scores by aligning even more closely with the directionality of benign updates than the benign clients themselves.
  • Credit Penalty for Filtered Clients. This mechanism addresses the risk of adaptive adversaries regaining trust too easily after being filtered. Specifically, clients removed during the Normalization Judgment step are presumed to have produced non-normalized or low-quality gradients and thus receive a credit decay to reduce their future influence. In contrast, clients filtered via the Outlier Suppression mechanism may have presented gradients that appear highly confident yet abnormally extreme. These gradients often arise from adversarial strategies that construct malicious updates by scaling deviations from the benign gradient mean, making them closely aligned with benign directionality and capable of evading baseline checks. Since such clients may initially appear trustworthy, a reverse penalty (i.e., increasing their credit score using $\gamma_{2}>1$) makes them more likely to be flagged again in subsequent outlier filtering, reinforcing detection consistency across rounds.